# Are We Learning the Right Features? A Framework for Evaluating DL-Based Software Vulnerability Detection Solutions

### v3

@@ Abstract
I Introduction
@@
VIPer, a perturbation-based framework111https://anonymous.4open.science/r/viper-4FD3/README.md to gauge the robustness of vulnerability detectors;
@@ II-B Identifying Spurious Features (SF)
To this end, we focused on the assumptions made by graph-based detectors [12, 13] regarding features that may contribute to a vulnerability. For instance, one common assumption is that a vulnerability is defined strictly by the set of nodes in the graph of the vulnerable sample. However, we observe that the same vulnerability would still exist if a mock node (e.g., a print("") statement) is added to the graph. Another common assumption is that the set of edges strictly defines the vulnerability in the graph of the vulnerable sample. However, a mock edge (e.g., if (5!=5) return;) can be added to the graph without affecting its vulnerability. The idea of introducing changes that should not impact vulnerability to the sets of nodes and edges was inspired by the existing literature that discussed how graph neural networks learn spurious correlations between sets of nodes and edges [26]. Our examination of the SARD dataset revealed that the above two assumptions in particular do not always hold and can lead to spurious correlations that impact detectors’ predictions. We demonstrate the extent of this impact in Section V.
@@ II-C Developing the Taxonomy
To capture and study the dichotomy between VFs and SFs, and to provide a comprehensive understanding of their characteristics, we classified them into a taxonomy of Code Features. The taxonomy is shown in Fig. 2. While it has proven indispensable as the basis for VIPer’s perturbations and accompanying analyses discussed below, the taxonomy is not intended to be comprehensive. We expect that follow-on work will add further categories to the taxonomy, which will in turn be encoded as further VIPer rules.
We employed the hierarchy-based taxonomy guidelines outlined in [27] to structure our taxonomy, with Code Features as the top-level class that encompasses all sub-classes via an ‘is-a’ relationship. We then partitioned Code Features into the two mutually exclusive sub-classes Vulnerability Features (VFs) and Spurious Features (SFs).
All 10 identified VFs from Table LABEL:tab:top10cwes are classified under the Vulnerability Features sub-class, which is subdivided into four further sub-classes based on feature characteristics:
@@
We categorized the Spurious Features based on the two primary approaches for Deep Learning (DL)-based vulnerability detection:
@@
Graph-based SFs involve relationships between nodes and edges that exhibit spurious correlations but do not inherently contribute to vulnerabilities. We categorize these relationships into (a) Node Set and (b) Edge Set.
The transition from individual structured annotations in the SARD dataset to the development of the taxonomy was carried out iteratively. We analyzed the annotated descriptions of each sample containing VFs. During each iteration, an author suggested a category and potential sub-categories, which were then discussed and refined. This process continued until all authors reached a consensus. Additionally, we incorporated categories for SFs by consulting relevant literature and applying the same iterative approach.
@@ III Approach
Our approach to Vulnerability Identification via Perturbations, VIPer, comprises three phases, as illustrated in Fig. 3: (1) Feature Detection uses CPG rules to identify the presence of VFs in samples and annotate them as candidates for perturbation. Since SFs are inherent properties of code, this phase does not pertain to them. (2) Targeted Perturbation systematically applies FPPs (alterations in code with no change to a feature) and FEPs (alteration in code that eliminates the feature) to the tagged samples, generating a perturbed dataset for each of the 10 VFs. For SFs, VIPer only generates FPPs because SFs are inherent properties of code that cannot be eliminated. We used the structured annotations in the SARD dataset as ground truth for evaluating the correctness of the CPG rules to detect features. We manually compared the CPG rule-detected features in 363 out of the 6,525 files ( confidence with a margin of error) with the annotations in SARD. When devising the FPPs and FEPs, we ensured that the perturbation targeted the property mentioned in the annotations and later manually analyzed the same 363 files to determine if the FPPs retained the vulnerability and the FEPs removed the vulnerability. (3) Solution Evaluation assesses the detectors’ responses to the perturbed datasets, examining how perturbations affect the predictive performance of the detectors and their ability to react satisfactorily. A satisfactory reaction is defined as no prediction change for FPPs and a prediction change for FEPs. With this information, VIPer quantifies the influence of these features on the detectors’ predictions, providing insights into their robustness and reliability. The following subsections will detail each phase.
@@ III-A1 CPG extraction
A CPG is formally defined as a graph , where is the set of all nodes and the set of all edges in the CPG, and is a property function that sets or retrieves the property of a node or edge. is the key that denotes which property would be retrieved or set, and is the value for property . To facilitate our implementation of VIPer, we extended the property function ’s capabilities to retrieve and set additional properties of nodes and edges. Due to space limitation, we list these additional property keys and values (as acronyms) used by property function in the online appendix.222https://sites.google.com/view/viper-framework
@@ III-A2 Pattern matching
To assess if a CPG satisfies a rule, we develop corresponding detection algorithms for each of the rules listed in Table II. Due to the page limitation, we will only describe the detection algorithm for Rule ID: 2.1 for VF Incorrect Calculation of Buffer Size which is the CPG rule that the sample in Listing 1 satisfies. We provide the algorithms for the remaining detection algorithms in the online appendix22footnotemark: 2.
@@
2.1 | Incorrect Calculation of Buffer Size (IBS) | CPG constructed from the sample must have a node representing a buffer write function call that writes number of bytes to a buffer with a defined length of that is smaller than (i.e., ). |
2.2 | Buffer Access Using Size of Source Buffer (BSB) | CPG constructed from the sample must have a node representing a buffer write function call that writes number of bytes from a source buffer with a defined length of equaling to a destination buffer with a defined length of that is smaller than (i.e., ). |
2.3 | Off-by-one Error (OE) | CPG constructed from the sample must have a node representing a buffer write function call that writes number of bytes to a buffer with a defined length of that is smaller than by exactly one. (i.e., ). |
2.4 | Buffer Over-read (BO) | CPG constructed from the sample must have a node representing a buffer copy function call that reads number of bytes from a buffer with a defined length of that is smaller than (i.e., ). |
2.5 | Double-Free (DF) | CPG constructed from the sample must have two nodes and who call free on the same buffer and there exists no node between and that uses an allocation function (e.g., malloc) on . |
2.6 | Use-After-Free (UAF) | CPG constructed from the sample must have two nodes and where node uses a buffer after node already deallocates buffer . |
2.7 | Buffer Underwrite (BUW) | CPG constructed from the sample must have a node that writes to a buffer using an index value where is never checked to ensure that it does not hold a negative value. |
2.8 | Buffer Under-read (BUR) | CPG constructed from the sample must have a node that reads from a buffer using an index value where is never checked to ensure that it is not a negative number. |
2.9 | Read API (RA) | CPG constructed from the sample must have a node representing a function call to a sensitive Read API (e.g., fgets) where the location of node is a vulnerable line in the sample. |
2.10 | Write API (WA) | CPG constructed from the sample must have a node representing a function call to a sensitive Write API (e.g., memcpy) where the location of node is a vulnerable line in the sample. |
@@ III-A3 Sample annotation
Once a feature is detected in a sample, VIPer annotates the sample with the detected feature, relevant line numbers, and variable names, which differ across the 10 VFs. For example, when VIPer detects Incorrect Calculation of Buffer Size in a sample (like in Listing 1), it first lists the name of the detected feature as Incorrect Calculation of Buffer Size and lists the line number that defines (line 4 in Listing 1) and the line number where the value of is defined (line 7 in Listing 1) and the variable name used for the destination buffer (data for Listing1). We provide the annotations used for the remaining VFs in the online appendix.22footnotemark: 2 After completing the sample annotations, VIPer creates a dataset comprising only the annotated samples. This annotated dataset is utilized by the subsequent phase of VIPer, Targeted Perturbation, which we will describe next.
@@ III-B Targeted Perturbation
Fig. 3 provides a high-level overview of how VIPer generates perturbations. The input for this phase is the annotated dataset generated by the previous phase, Feature Detection. From this dataset, VIPer extracts the three essential elements required for perturbation: (1) the original sample in the dataset where the VF was detected, (2) the name of the detected VF and (3) the relevant line numbers and variable names for generating perturbations. Based on the detected feature’s name, VIPer selects one of 10 tailored perturbation generation algorithm sets. Each set involves generating one or more FPPs and FEPs. Due to page limitations, we will only elaborate on the perturbation generation algorithm set for Incorrect Calculation of Buffer Size. The remaining nine algorithm sets are provided in the online appendix22footnotemark: 2. Recall the CPG rule for VF Incorrect Calculation of Buffer Size is . The algorithms used for generating FPPs and FEPs for this CPG rule are given as follows:
@@ III-B2 FEP generation
In contrast, for graph-based SFs, we develop new perturbations. Importantly, since modifications to SFs do not impact the sample’s vulnerability ground truth, all generated perturbations for SFs are FPPs. The goal for SF perturbations is to modify the nodes and edges in a sample with minimal possible change, without affecting the vulnerability ground truth. Therefore, for Node Set, the corresponding perturbation is inserting a printf(""); statement at the start of each function and for Edge Set, the corresponding perturbation is inserting if(0==1) return; at the start of each function. Note that these perturbations are generic and are applied to all samples in the dataset.
@@ IV-B Classifying Evaluation Results
For the purpose of our analysis and discussion, specifically RQ3, we classify detector responses into one of four -- combinations: high--high (HH), high--low (HL), low--high (LH), and low--low (LL). We consider the satisfaction rate high if the value is within (chosen value ) of the average satisfaction rate for a perturbation type targeting feature . This ensures that detectors performing close to the average are still recognized as effective.333While our analysis would be carried out the same way regardless of the selected value , we selected this value based on established conventions, as an of is frequently used in surveys, such as the American Statistical Association (https://www.amstat.org/docs/default-source/amstat-documents/pol-seeingthroughstats_spring_2013.pdf) and the American Community Survey’s methodology for population and housing estimates (https://www.census.gov/programs-surveys/acs). Detectors in the HH category exhibit a desired understanding of feature since they retain predictions for FPPs and change predictions for FEPs. These detectors will not experience significant drop in precision or recall in the presence of perturbations. A detector in the HL category exhibits an understanding of feature , but that understanding does not align with the ground-truth characteristics of (i.e., the relevant CPG rule). These detectors will experience a noticeable drop in precision in the presence of perturbations. A detector in the LH category changes its predictions given any perturbation involving feature . Such detectors will experience a significant drop in recall. Lastly, detectors in the LL category will exhibit a significant drop in both precision and recall in the presence of perturbations.
Researchers can use VIPer’s results to assess the effectiveness of vulnerability detectors and decide which option is best suited for their purpose. Ideally, one would select a detector falling under the HH category. If none of the available detectors fall under that category, an engineer has to assess whether a detector within one of the other categories is suitable for their tasks. For example, if a task prioritizes recall over precision, detectors falling under the HL category may be considered.
@@ IV-C Evaluation Subjects
We investigated five state-of-the-art vulnerability detectors to evaluate whether they learn from VFs or SFs: DeepWukong [12], Reveal[13], DeepDFA [20], LineVul [14], and SySeVR [11].
@@ IV-D Evaluation Dataset
To assess the five evaluation subjects, VIPer uses the largest vulnerability dataset, SARD, containing production, synthetic, and academic programs (also known as test cases). We utilized the curated version of SARD from SySeVR as it is the largest dataset used among all five vulnerability detectors.444https://github.com/SySeVR/SySeVR/tree/master/Program%20data/SARD The dataset includes 22,080 C/C++ files. The vulnerable programs in SARD provide precise locations of each vulnerability, enabling effective analysis. In total, the dataset had 366,419 C/C++ functions that we used for training the five evaluation subjects.
@@ V-A RQ1: Analysis of Accuracy
For VFs that belong to Deallocated Buffer Use vulnerability sub-class, we observe that all graph-based detectors (DeepWukong, ReVeal, and DeepDFA) noticeably outperform the token-based detectors (LineVul and SySeVR). DeepWukong achieves the smallest drop in precision ( on average) and a slight increase in recall ( on average), ighest drop in precision (up to ). Since both token-based approaches demonstrate relatively poor performance (i.e., over drop in precision on average for LineVul and similar drop in recall for SySeVR), this could be due to the fact that Deallocated Buffer Use vulnerabilities are characterized mainly by the control flow of a program [25]. Since graph-based detectors, by design, are able to better capture context information from graphs, they are expected to outperform token-based approaches. However, DeepDFA only focuses on the dataflow of a function which may explain its drop in precision. For VFs that belong to Numeric Range Comparison vulnerabilities, LineVul shows the highest drop in precision () among the five detectors.
Graph-based | Token-based | |||||
Feature Name | DeepWukong | ReVeal | DeepDFA | LineVul | SySeVR | |
@@
For the SF Code Formatting, we observe that both SySeVR and LineVul are reasonably resilient with the drop in precision and recall never exceeding . For the SF Identifier Name, LineVul shows a significant drop in precision with over , while SySeVR’s precision drops only by . This is likely because SySeVR preprocesses raw code tokens (e.g., symbolizing code elements to prevent learning from custom naming conventions) while LineVul converts the raw code tokens directly into vectors thus exposing itself to this SF. For two SFs that belong to the graph-based sub-class, we observe that DeepWukong only suffers a noticeable drop in recall for the SF Edge Set (), while ReVeal’s recall drops for both SFs. This reduction in recall in both cases suggests that graph-based approaches tend to generate false negatives whenever encountering a behavior-preserving perturbation, such as introducing a mock edge to the graph. However, in case of Node Set perturbations, DeepWukong is less influenced, possibly due to the fact that it only considers control and data dependence edges from the CPG; since Node Set perturbations do not introduce any of these dependencies, they do not influence DeepWukong’s predictions. In contrast, ReVeal considers all the nodes from CPG, and is thus inherently exposed to Node Set perturbations. The precision and recall drops for DeepDFA are negligible (). Similarly to the above discussion of DeepWukong, this is likely due to the fact that DeepDFA only works on dataflow information: since the SF perturbations introduce no changes to a program’s dataflow, DeepDFA is able to effectively filter them out during inference.

### v5

@@ Abstract
Index Terms:
vulnerability detection, deep learning, software security, explainable AII Introduction
@@
VIPer, a perturbation-based framework [21] to gauge the robustness of vulnerability detectors;
@@ II-B Identifying Spurious Features (SF)
To this end, we focused on the assumptions made by graph-based detectors [12, 13] regarding features that may contribute to a vulnerability. For instance, one common assumption is that a vulnerability is defined strictly by the set of nodes in the graph of the vulnerable sample. However, we observe that the same vulnerability would still exist if a mock node (e.g., a printf("Benign") statement) is added to the graph. Another common assumption is that the set of edges strictly defines the vulnerability in the graph of the vulnerable sample. However, a mock edge (e.g., if (5!=5) return;) can be added to the graph without affecting its vulnerability. The idea of introducing changes that should not impact vulnerability to the sets of nodes and edges was inspired by the existing literature that discussed how graph neural networks learn spurious correlations between sets of nodes and edges [27]. Our examination of the SARD dataset revealed that the above two assumptions in particular do not always hold and can lead to spurious correlations that impact detectors’ predictions. We demonstrate the extent of this impact in Section V.
@@ II-C Developing the Taxonomy
To capture the dichotomy between VFs and SFs, and to provide a comprehensive understanding of their characteristics, we classified them into a taxonomy of Code Features. Our hierarchy-based taxonomy [28] is shown in Fig. 2. The taxonomy is not intended to be comprehensive. We expect that follow-on work will add further categories to the taxonomy, which will in turn be encoded as further VIPer rules.
We partitioned Code Features into two mutually exclusive sub-classes Vulnerability Features (VFs) and Spurious Features (SFs). All 10 identified VFs from Table LABEL:tab:top10cwes are classified under the Vulnerability Features sub-class, which is subdivided into four further sub-classes based on feature characteristics:
@@
We categorized the Spurious Features based on the two primary approaches for DL-based vulnerability detection:
@@
Graph-based SFs are features of nodes and their relationships (edges) that exhibit spurious correlations but do not inherently contribute to vulnerabilities. We categorize these features into (a) Node Set and (b) Edge Set. Node Set refers to the presence of one or more single-line instructions that have no impact on the vulnerability. Specific instances of these SFs include (i) Print instructions without variables, i.e., ‘‘empty’’ print statements in the code, and (ii) Unused variable declarations, i.e., variables that are not used after declaration. Edge Set refers to the presence of multi-line instructions defining data or control dependencies that do not affect vulnerability. Specific instances of Edge Set vulnerabilities are (i) Unreachable conditional operation without variables, i.e., a condition that will always be false and may change the control dependency, and (ii) Dependency-altering uses of variables, i.e., usage of variables that may alter dataflow or control dependency but that do not contribute to vulnerability. Note that the goal of the study reported in this paper is to observe the impact of SF categories that require minimal change. For this reason, we exclude the SFs that are greyed out in Fig. 2 from this study.
The transition from individual structured annotations in the SARD dataset to the development of the taxonomy was carried out iteratively. We analyzed the annotated descriptions of each sample containing VFs. During each iteration, an author suggested a category and potential sub-categories, which were then discussed and refined. This process continued until all authors reached a consensus. Additionally, we incorporated categories for SFs by consulting relevant literature [12, 13, 27] and applying the same iterative approach.
@@ III Approach
Our approach to Vulnerability Identification via Perturbations, VIPer, comprises three phases, as illustrated in Fig. 3: (1) Feature Detection uses CPG rules to identify the presence of VFs in samples and annotate them as candidates for perturbation. Since SFs are inherent properties of code, this phase does not pertain to them. (2) Targeted Perturbation systematically applies FPPs (alterations in code with no change to a feature) and FEPs (alteration in code that eliminates the feature) to the tagged samples, generating a perturbed dataset for each of the 10 VFs. For SFs, VIPer only generates FPPs because SFs are inherent properties of code that cannot be eliminated. We used the structured annotations in the SARD dataset as ground truth for evaluating the correctness of the CPG rules to detect features. We manually compared the CPG rule-detected features in 363 out of the 6,525 files ( confidence with a margin of error) with the annotations in SARD. When devising the FPPs and FEPs, we ensured that the perturbation targeted the property mentioned in the annotations and later manually analyzed the same 363 files to determine if the FPPs retained the vulnerability and the FEPs removed the vulnerability. (3) Solution Evaluation assesses the detectors’ responses to the perturbed datasets, examining how perturbations affect the predictive performance of the detectors and their ability to react satisfactorily. A satisfactory reaction is defined as no prediction change for FPPs and a prediction change for FEPs. With this information, VIPer quantifies the influence of these features on the detectors’ predictions. The following subsections detail each phase.
@@ III-A1 CPG extraction
A CPG is formally defined as a graph , where is the set of all nodes and the set of all edges in the CPG, and is a function that sets or retrieves the property of a node or edge. is the key that denotes which property would be retrieved or set, and is the value for property . To facilitate our implementation of VIPer, we extended the property function ’s capabilities to retrieve and set additional properties of nodes and edges. Due to space limitation, we list these additional property keys and values (as acronyms) used by property function in the online appendix [31].
@@ III-A2 Pattern matching
To assess if a CPG satisfies a rule, we develop corresponding detection algorithms for each of the rules listed in Table II. Due to the page limitation, we will only describe the detection algorithm for Rule ID: 2.1 for VF Incorrect Calculation of Buffer Size which is the CPG rule that the sample in Listing 1 satisfies. We provide the algorithms for the remaining detection algorithms in the online appendix [31].
@@
| 2.1 | Incorrect Calculation of Buffer Size (IBS) | CPG constructed from the sample must have a node representing a buffer write function call that writes number of bytes to a buffer with a defined length of that is smaller than (i.e., ). |
| 2.2 | Buffer Access Using Size of Source Buffer (BSB) | CPG constructed from the sample must have a node representing a buffer write function call that writes number of bytes from a source buffer with a defined length of equaling to a destination buffer with a defined length of that is smaller than (i.e., ). |
| 2.3 | Off-by-one Error (OE) | CPG constructed from the sample must have a node representing a buffer write function call that writes number of bytes to a buffer with a defined length of that is smaller than by exactly one. (i.e., ). |
| 2.4 | Buffer Over-read (BO) | CPG constructed from the sample must have a node representing a buffer copy function call that reads number of bytes from a buffer with a defined length of that is smaller than (i.e., ). |
| 2.5 | Double-Free (DF) | CPG constructed from the sample must have two nodes and who call free on the same buffer and there exists no node between and that uses an allocation function (e.g., malloc) on . |
| 2.6 | Use-After-Free (UAF) | CPG constructed from the sample must have two nodes and where node uses a buffer after node already deallocates buffer . |
| 2.7 | Buffer Underwrite (BUW) | CPG constructed from the sample must have a node that writes to a buffer using an index value where is never checked to ensure that it does not hold a negative value. |
| 2.8 | Buffer Under-read (BUR) | CPG constructed from the sample must have a node that reads from a buffer using an index value where is never checked to ensure that it is not a negative number. |
| 2.9 | Read API (RA) | CPG constructed from the sample must have a node representing a function call to a sensitive Read API (e.g., fgets) where the location of node is a vulnerable line in the sample. |
| 2.10 | Write API (WA) | CPG constructed from the sample must have a node representing a function call to a sensitive Write API (e.g., memcpy) where the location of node is a vulnerable line in the sample. |
@@ III-A3 Sample annotation
Once a feature is detected in a sample, VIPer annotates the sample with the detected feature, relevant line numbers, and variable names, which differ across the 10 VFs. For example, when VIPer detects Incorrect Calculation of Buffer Size in a sample (like in Listing 1), it first lists the name of the detected feature as Incorrect Calculation of Buffer Size and lists the line number that defines (line 4 in Listing 1) and the line number where the value of is defined (line 7 in Listing 1) and the variable name used for the destination buffer (data for Listing1). We provide the annotations used for the remaining VFs in the online appendix [31]. After completing the sample annotations, VIPer creates a dataset comprising only the annotated samples. This annotated dataset is utilized by the subsequent phase of VIPer, Targeted Perturbation, which we will describe next.
@@ III-B Targeted Perturbation
Fig. 3 provides a high-level overview of how VIPer generates perturbations. The input for this phase is the annotated dataset generated by the previous phase, Feature Detection. From this dataset, VIPer extracts the three essential elements required for perturbation: (1) the original sample in the dataset where the VF was detected, (2) the name of the detected VF and (3) the relevant line numbers and variable names for generating perturbations. Based on the detected feature’s name, VIPer selects one of 10 tailored perturbation generation algorithm sets. Each set involves generating one or more FPPs and FEPs. Due to page limitations, we will only elaborate on the perturbation generation algorithm set for Incorrect Calculation of Buffer Size. The remaining nine algorithm sets are provided in the online appendix [31]. Recall the CPG rule for VF Incorrect Calculation of Buffer Size is . The algorithms used for generating FPPs and FEPs for this CPG rule are given as follows:
@@ III-B2 FEP generation
In contrast, for graph-based SFs, we develop new perturbations. Importantly, since modifications to SFs do not impact the sample’s vulnerability ground truth, all generated perturbations for SFs are FPPs. The goal for SF perturbations is to modify the nodes and edges in a sample with minimal possible change, without affecting the vulnerability ground truth. Therefore, for Node Set, the corresponding perturbation is inserting a printf(""); statement at the start of each function (i.e., the Print instructions without variables SF) and for Edge Set, the corresponding perturbation is inserting if(0==1) return; at the start of each function (i.e., the Unreachable conditional operation without variables SF). 111For better readability, we are referring to the SFs Print instructions without variables and Unreachable conditional operation without variables as Node Set and Edge set, respectively, for the rest of the paper. Note that these perturbations are generic and are applied to all samples in the dataset.
@@ IV-B Classifying Evaluation Results
For the purpose of our analysis and discussion, specifically RQ3, we classify detector responses into one of four -- combinations: high--high (HH), high--low (HL), low--high (LH), and low--low (LL). We consider the satisfaction rate high if the value is within (chosen value ) of the average satisfaction rate for a perturbation type targeting feature . This ensures that detectors performing close to the average are still recognized as effective.222 While our analysis would be carried out the same way regardless of the selected value , we selected this value based on established conventions, as an of is frequently used in surveys, such as the American Statistical Association [33] and the American Community Survey’s methodology for population and housing estimates [34]. Detectors in the HH category exhibit a desired understanding of feature since they retain predictions for FPPs and change predictions for FEPs. These detectors will not experience significant drop in precision or recall in the presence of perturbations. A detector in the HL category exhibits an understanding of feature , but that understanding does not align with the ground-truth characteristics of (i.e., the relevant CPG rule). These detectors will experience a noticeable drop in precision in the presence of perturbations. A detector in the LH category changes its predictions given any perturbation involving feature . Such detectors will experience a significant drop in recall. Lastly, detectors in the LL category will exhibit a significant drop in both precision and recall in the presence of perturbations.
Researchers and engineers can use VIPer’s results to assess the vulnerability detectors and decide which option is best suited for their purpose. Ideally, one would always prefer a detector falling under the HH category. If none of the available detectors fall under that category, their users will have to assess whether a detector that falls within one of the other categories is suitable for their tasks. Generally, vulnerability detection tasks prioritize recall over precision, i.e., tolerating false alarms while fixing as many vulnerabilities as possible. In such cases, the engineer may prefer detectors falling under the HL category. However, there may be certain tasks where precision is preferred over recall. An example is as a large and stable system in which vulnerable code is less prevalent and the cost of handling false alarms is significant. In that case, the LH category may be considered.
@@ IV-C Evaluation Subjects
We investigated whether five state-of-the-art vulnerability detectors learn from VFs or SFs: DeepWukong [12], Reveal[13], DeepDFA [20], LineVul [14], and SySeVR [11].
@@ IV-D Evaluation Dataset
To assess the five evaluation subjects, VIPer uses the largest vulnerability dataset, SARD, containing production, synthetic, and academic programs (a.k.a. test cases). We utilized the curated version of SARD from SySeVR as it is the largest dataset used among all five vulnerability detectors [36]. The dataset includes 22,080 C/C++ files. The vulnerable programs in SARD provide precise locations of each vulnerability, enabling effective analysis. In total, the dataset had 366,419 C/C++ functions that we used for training the five evaluation subjects.
@@ V-A RQ1: Analysis of Accuracy
For VFs that belong to Deallocated Buffer Use vulnerability sub-class, we observe that all graph-based detectors (DeepWukong, ReVeal, and DeepDFA) noticeably outperform the token-based detectors (LineVul and SySeVR). DeepWukong achieves the smallest drop in precision ( on average) and a slight increase in recall ( on average), while DeepDFA exhibits the highest drop in precision (up to ). Since both token-based approaches demonstrate relatively poor performance (i.e., over drop in precision on average for LineVul and similar drop in recall for SySeVR), this could be due to the fact that Deallocated Buffer Use vulnerabilities are characterized mainly by the control flow of a program [26]. Since graph-based detectors, by design, are able to better capture context information from graphs, they are expected to outperform token-based approaches. However, DeepDFA only focuses on the dataflow of a function which may explain its drop in precision. For VFs that belong to Numeric Range Comparison vulnerabilities, LineVul shows the highest drop in precision () among the five detectors.
| Graph-based | Token-based | ||||||||||
| Feature Name | DeepWukong | ReVeal | DeepDFA | LineVul | SySeVR | ||||||
@@
For the SF Code Formatting, we observe that both SySeVR and LineVul are reasonably resilient with the drop in precision and recall never exceeding . For the SF Identifier Name, LineVul shows a significant drop in precision with over , while SySeVR’s precision drops only by . This is likely because SySeVR preprocesses raw code tokens (e.g., symbolizing code elements to prevent learning from custom naming conventions) while LineVul converts the raw code tokens directly into vectors thus exposing itself to this SF. For two SFs that belong to the graph-based sub-class, we observe that DeepWukong only suffers a noticeable drop in recall for the SF Edge Set (), while ReVeal’s recall drops for both SFs. This reduction in recall in both cases suggests that graph-based approaches tend to generate false negatives whenever encountering a behavior-preserving perturbation, such as introducing a mock edge to the graph. However, in case of Node Set perturbations, DeepWukong is less influenced, possibly due to the fact that it only considers control and data dependence edges from the CPG; since Node Set perturbations do not introduce any of these dependencies, they do not influence DeepWukong’s predictions. In contrast, ReVeal considers all the nodes from CPG, and is thus inherently exposed to Node Set perturbations. The precision and recall drops for DeepDFA are negligible (). Similarly to the above discussion of DeepWukong, this is due to the fact that DeepDFA only works on dataflow information and SF perturbations introduce no changes to a program’s dataflow. At the same time, it may be worth investigating whether DeepDFA retains the same robustness against SFs that impact dataflow information, e.g., the use of intermediate variables during a mathematical calculation.